• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lingyan

    Log entry time 00:07:54 on June 4,2001

    Entry number 61672

    keyword=swing shift summary

    16:00 worked on MCC magnets, harp scan ioc
    20:40 beam available, did harp scan to center the beam on the target
    21:56 moved target to pick-up coil and did NMR measurement
    23:00 moved target to empty and increased current to 10uA, for compton measurment.
    MCC said the maximum current they can deliver is 10uA.


    Beam energy: 1.2GeV,
    Momentum settng: 1.1903GeV/c, 19.515 deg(Left), 22.347 deg(Right)


    target...left_run#...start/stop...right_run#...start/stop...comments
    cosmic...1315......16:37/20:03.....20316......16:46/20.02
    BeO.......1316......21:23/22:47.....20317.......21:23/22:47...beam position adjustment
    empty....1318.......23:08/23:46
    empty....1319.......23:48/-...........20319........23:08/-..........10uA, lots of beam trips

    pass to next shift